Project

General

Profile

RunPlan20150611 » History » Revision 3

Revision 2 (Blackmore, Victoria, 11 June 2015 18:35) → Revision 3/6 (Blackmore, Victoria, 11 June 2015 18:37)

h1. June 11th, 2015 (Evening Run) 

 *Aim:* Calibrating TOFs, testing Run Control, testing Online Reconstruction 


 h3. Daily schedule 

 | June    | 11th |    
 |Time |    9pm - 5am | 
 | MOM         | V. Blackmore     |   
 | Lead shifter | S. Middleton |  
 | Shifter 2 |    J. Nugent    |  
 | BLOC        | A. Dobbs     |  
 | SOC         | A. Dobbs     |  
 | Experimenter | A. Dobbs, P. Hanlet    |  
 | Physics Shifter | P. Franchini | 
 |    | | |  
 | Beam?    | Yes    |  
 | Additional Personnel | | |   


 h2. Shift plan: 

 * 7:30pm 
 ** Shifters + BLOC arrive 
 ** Hall and beam line to be closed and prepared for beam 
 * 9pm 
 ** E. Overton to remove tracker veto from DAQ before TOF calibration begins 
 ** Check TOF1 response channel-by-channel  
 ** Operate experiment via Run Control, take data through to TOF2/KL/EMR to calibrate TOFs and test Online Reconstruction 
 ** Settings as A. Dobbs requires 


 h2. Data Requested 

 * TOF1 calibration 
 * TOF0 and TOF2 calibration time permitting No data requested 


 h2. Beamline Configurations 

 * We expect to run with 2V beamloss and the beam bump active. 
 * DS is currently off 
 * Beam settings via Run Control 
 ** TOF1 calibration, short runs for Run Control testing 
 ** TOF1 calibration, long 100k triggers run -- aim for 1000 triggers/pixel 
 ** TOF0, TOF2 calibration, long 100k triggers runs (time permitting) 


 h2. Phone details for 6pm daily meeting 

 Phone access :    ReadyTalk number  

 US 866-740-1260 

 UK 0800-496-0576 

 Swiss 0800-70-53-51 

 Italy 800182592 

 Access Code: 2155402